# The Analysis of Willmore Surfaces and its Generalizations in Higher Dimensions

### v1

@@ IV.4.3 Approximation results
$\begin{aligned} &\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)-\big(d\pi\circ d(f\circ\phi_{j}^{-1})\big)_{z^{\prime}}(v)\big|\\[8.53581pt] &\leq\big|\nabla\pi_{(F_{\varepsilon}\circ\phi_{j}^{-1})(z)}\big(\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})_{z}-\nabla(f\circ\phi_{j}^{-1})_{z^{\prime}}\big)(v)\big|\\[7.96677pt] &\quad\quad+\big|\big(\nabla\pi_{(F_{\varepsilon}\circ\phi_{j}^{-1})(z)}-\nabla\pi_{(f\circ\phi_{j}^{-1})(z^{\prime})}\big)\nabla(f\circ\phi_{j}^{-1})_{z^{\prime}}(v)\big|\\[8.53581pt] &\leq\big|\nabla\pi\big((F_{\varepsilon}\circ\phi_{j}^{-1})(z)\big)\big|\ \big|\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-\nabla(f\circ\phi_{j}^{-1})(z^{\prime})\big|\\[9.95845pt] &\quad\quad+\text{Lip}(\nabla\pi|_{\overline{W^{\prime}}})\ \big|(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z^{\prime})\big|\ \big|\nabla(f\circ\phi_{j}^{-1})(z^{\prime})\big|.\end{aligned}$ (IV.80)
Combining (IV.76)–(LABEL:c0dis) and (LABEL:dpitri), we obtain
$\begin{aligned} &\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)-\big(d\pi\circ d(f\circ\phi_{j}^{-1})\big)_{z^{\prime}}(v)\big|\\[7.11317pt] &\leq(2\alpha)^{-1}C_{j}\Big(\|\nabla\pi\|_{C(\overline{W^{\prime}})}+\text{Lip}(\nabla\pi|_{\overline{W^{\prime}}})\,\|\nabla(f\circ\phi_{j}^{-1})\|_{L^{\infty}(\phi_{j}(V_{j}\cap\overline{U}))}\Big)\\[7.11317pt] &\leq 2^{-1}C_{j}.\end{aligned}$ (IV.81)
By (IV.75), (LABEL:dpi>), (LABEL:<c_j) we finally conclude that for any $z\in\phi_{j}(\operatorname{supp}(\xi_{j})\cap\overline{U})$, $0<\varepsilon<\varepsilon_{0}$, $v\in\mathbb{S}^{n-1}$, it holds
@@
The estimate (IV.72) is obtained since $\overline{U}\subset\bigcup_{j=1}^{s}\operatorname{supp}(\xi_{j})$. ∎
@@
$\Lambda^{-1}|X|_{\tilde{g}}\leq|d(\vec{\Phi}_{k})(X)|\leq\Lambda|X|_{\tilde{g}},\qquad\text{for all }X\in T\Sigma\,\text{ and all }\,k\in\mathbb{N}.$ (IV.82)
@@
Now it remains to show that $\alpha_{k}\rightarrow\alpha$ in $C^{0}(\Sigma)$. Let $p\in\Sigma$. Since $h_{k}^{\prime}\rightarrow h$ in $C^{\infty}(\Sigma)$, by a similar argument as in Theorem IV.17 (replacing $\vec{e}_{i,k}$ by orthonormal frames on $T\Sigma$, and $d\vec{e}_{i,k}$ by covariant derivatives of these vector fields with respect to $h^{\prime}_{k}$), there exists a neighborhood $U$ of $p$ and $C^{\infty}$ diffeomorphisms $\varphi_{k}$ from $U$ to $\varphi_{k}(U)\subset\mathbb{R}^{2}$ such that for any $k\in\mathbb{N}\cup\{\infty\}$, it holds that $\varphi_{k}(p)=0$, $D^{2}\Subset\varphi_{k}(U)$, and $\varphi_{k}^{*}g_{\text{std}}$ is conformal to $h_{k}^{\prime}$ (write $h_{\infty}^{\prime}\coloneqq h$). Moreover, we have $\varphi_{k}\rightarrow\varphi_{\infty}$ in $C^{\infty}_{\textup{loc}}(U)$. Since $h_{k}^{\prime}$ is conformal to $g_{k}^{\prime}$, we also obtain that $\vec{\Phi}_{k}^{\prime}\circ\varphi_{k}^{-1}$ is conformal. Write
@@
Then $\tilde{\lambda}_{k}\rightarrow\tilde{\lambda}_{\infty}$ in $C^{\infty}(\overline{D^{2}})$ since $\varphi_{k}^{-1}\rightarrow\varphi_{\infty}^{-1}$ in $C^{\infty}(\overline{D^{2}})$ and $h_{k}^{\prime}\rightarrow h_{\infty}^{\prime}=h$ in $C^{\infty}(\Sigma)$. Since $\{\lambda_{k}\}$ is bounded in $L^{\infty}(D^{2})$ by (IV.82) and Theorem IV.19 (v), and $\vec{\Phi}_{k}^{\prime}\circ\varphi_{k}^{-1}\rightarrow\vec{\Phi}\circ\varphi_{\infty}^{-1}$ in $W^{2,2}(D^{2},\mathbb{R}^{m})$, we have $\lambda_{k}\rightarrow\lambda_{\infty}$ in $W^{1,2}(D^{2})$. Now by applying the argument in Corollary III.11 to $\{\vec{\Phi}_{k}^{\prime}\circ\varphi_{k}^{-1}\}$, we have $\lambda_{k}\rightarrow\lambda_{\infty}$ in $W^{2,1}_{\textup{loc}}(D^{2})$. Since $\alpha_{k}=(\lambda_{k}-\tilde{\lambda}_{k})\circ\varphi_{k}$ on $U$, and $p$ is arbitrarily chosen on $\Sigma$, we conclude that $\alpha_{k}\rightarrow\alpha$ in $W^{2,1}(\Sigma)$, hence also in $C^{0}(\Sigma)$ (see [80, Theorem 3.3.4 & 3.3.10]). ∎
@@ V Willmore surfaces
Motivated by the generalization of Willmore surfaces in higher dimensions, we provide in this section, a new proof of the regularity of Willmore surfaces which does not involve the choice of conformal coordinates. In Section V.1, we write the Euler–Langrange equation of Willmore surfaces in divergence form. In Section V.2, we prove the regularity of Willmore surfaces.
@@ V.1 The Euler–Lagrange equation and conservation laws
Theorem V.1 (Noether’s theorem [154]).
@@
Thanks to the fact $\Delta_{g}\vec{\Phi}\perp\partial_{l}\vec{\Phi}$, we have
@@
Combing (V.4)–(V.7), we obtain the pointwise a.e. variation
@@
In particular, it holds that $d\vec{L}\in W^{-1,2}+L^{1}(D^{2})$, and by Corollary III.5, we have $\vec{L}\in L^{2}_{\text{loc}}(D^{2})$. The quantity $\vec{L}$ alone is not sufficient to deduce some additional regularity since $\vec{L}$ does not possess additional derivative a priori. Using the pointwise invariance of $H^{2}\,d\textup{vol}_{g}$ by dilations and rotations, we deduce some conservation laws related to $\vec{L}$ and $\vec{\Phi}$.
@@
Theorem V.4 ([168, Theorem 5.59]).
Let $\vec{\Phi}\in W^{2,2}_{\text{imm}}(D^{2},\mathbb{R}^{3})$ be a weak Willmore immersion. Then there exists $\vec{L}\in L^{2}_{\text{loc}}(D^{2},\mathbb{R}^{3})$ satisfying (V.15) [^8] Up to additive constants, this $\vec{L}$ is 1/2-1/2 of the one defined in [168, Theorem 5.59]. There is a typo in [168, Equation (5.214b)]: the ++ sign should be changed to -. and moreover, we have
@@
By differentiating the relations (V.19), we obtain a $\mathrm{div}-\operatorname{curl}$ system in $\vec{R}$ and $S$.
Theorem V.6 ([168, Corollary 5.61]).
@@
Although $H^{2}\,d\textup{vol}_{g}$ is not pointwise invariant by inversions, the Lagrangian $(H^{2}-K)\,d\textup{vol}_{g}$ is; see for instance [34]. In this section, we derive the associated conservation law (V.28), which is equivalent to (V.20c).
@@ V.2 Regularity of Willmore surfaces without conformal coordinates
By dilation and translation, it suffices to prove $\nabla S,\nabla\vec{R},\vec{H}\in M^{2,\gamma}(D_{1/2}(0))$. By definition of a weak immersion, the coefficients $a^{ij}\coloneqq g^{ij}(\det g)^{1/2}$ satisfies the uniform ellipticity condition (III.13). Fix a positive small number $\varepsilon_{0}$ which will be determined later. Since $\vec{\Phi}\in W^{2,2}(D^{2})$, we have $\nabla\vec{n}\in L^{2}(D^{2})$, hence there exists $r_{0}\in(0,\frac{1}{4})$ depending on $\epsilon_{0}$ and $\vec{\Phi}$ such that
@@
Let $p\in D_{1/2}(0)$, $r\in(0,r_{0})$. By Lemma III.3, we can define $\Psi_{S}\in W^{1,2}_{0}(D_{r}(p))$ to be the unique solutions of (see Section III.1)
@@
We can also define $\vec{\Psi}_{\vec{R}}\in W^{1,2}_{0}(D_{r}(p))$ to be the unique solutions of
@@
Fix a positive small number $\varepsilon_{0}$ which will be determined later. As in the proof of Theorem V.8, there exists $r_{0}\in(0,\frac{1}{4})$ depending on $\epsilon_{0}$ and $\vec{\Phi}$ such that
@@
Furthermore, by Riesz representation theorem for Hilbert spaces, there exists a unique $\vec{\Psi}_{2}\in W^{1,2}_{0}(D_{r}(p))$ satisfying the estimate $\|\nabla\vec{\Psi}_{2}\|_{L^{2}(D_{r}(p))}\leq C(\Lambda)\,\|H\|_{L^{2}(D_{r}(p))}$ and the system
@@
The Hölder continuity of $\vec{n}$ then follows from standard knowledge of the Morrey–Campanato spaces (see for instance [150, Theorem 3.5.2] and [2]). ∎
@@
Let $\Sigma$ be a $2$-dimensional closed smooth manifold, $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(\Sigma,\mathbb{R}^{3})$. Then we have $\vec{\Phi}$ is a graph near $\vec{\Phi}(p)$ for any $p\in\Sigma$, and $\vec{\Phi}\in C^{\infty}$ under the graph coordinates.
@@
By [145, Eq. (1.5)] and standard estimates on the Riesz potential [1, Proposition 3.2(ii)], we obtain $M^{1,\tau}(D^{2})\hookrightarrow W^{-1,p}(D^{2})$ for any $p\in(2,\frac{2-\tau}{1-\tau})$. Now since $_{g}\big(d\vec{n}\mathbin{\dot{\wedge}}d\vec{R}\big)\in M^{1,\tau}_{\textup{loc}}(D^{2})$, we have $_{g}\big(d\vec{n}\mathbin{\dot{\wedge}}d\vec{R}\big)\in W^{-1,p}_{\textup{loc}}(D^{2})$ for some $p>2$. Applying [31, Theorem 3.1] to the equation (V.43a) then yields $\nabla S\in L^{p}_{\textup{loc}}(D^{2})$. Similarly, we have $\nabla\vec{R}\in L^{p}_{\textup{loc}}(D^{2})$, and hence by (V.43c), $H\in L^{p}_{\textup{loc}}(D^{2})$. Moreover, since $d(2H*_{g}d\vec{\Phi})\in W^{-1,p}_{\textup{loc}}(D^{2})$, combining (V.43d) and (V.44) with a similar argument as above implies $\nabla\vec{n}\in L^{p}_{\textup{loc}}(D^{2})$. It follows that
@@
Now return to the equations (V.43a)–(V.43d). The right hand side of each equation is in $L^{q}_{\textup{loc}}(D^{2})$ for any $q<\infty$, and since $S,\vec{R},\vec{n},g\in W^{1,q}_{\textup{loc}}(D^{2})$, by [38, Theorem 4.1] and the proof of [66, Theorem 8.8], we obtain $S,\vec{R},\vec{n}\in W^{2,q}_{\textup{loc}}(D^{2})$, and then $\vec{\Phi}\in W^{3,q}_{\textup{loc}}(D^{2})$… We finally get $\vec{\Phi}\in C^{\infty}$ under the graph coordinates. ∎

### v3

@@ IV.4.3 Approximation results
$\begin{aligned} &\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)-\big(d\pi\circ d(f\circ\phi_{j}^{-1})\big)_{z^{\prime}}(v)\big|\\[8.53581pt] &\leq\big|\nabla\pi_{(F_{\varepsilon}\circ\phi_{j}^{-1})(z)}\big(\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})_{z}-\nabla(f\circ\phi_{j}^{-1})_{z^{\prime}}\big)(v)\big|\\[7.96677pt] &\quad\quad+\big|\big(\nabla\pi_{(F_{\varepsilon}\circ\phi_{j}^{-1})(z)}-\nabla\pi_{(f\circ\phi_{j}^{-1})(z^{\prime})}\big)\nabla(f\circ\phi_{j}^{-1})_{z^{\prime}}(v)\big|\\[8.53581pt] &\leq\big|\nabla\pi\big((F_{\varepsilon}\circ\phi_{j}^{-1})(z)\big)\big|\ \big|\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-\nabla(f\circ\phi_{j}^{-1})(z^{\prime})\big|\\[9.95845pt] &\quad\quad+\text{Lip}(\nabla\pi|_{\overline{W^{\prime}}})\ \big|(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z^{\prime})\big|\ \big|\nabla(f\circ\phi_{j}^{-1})(z^{\prime})\big|.\end{aligned}$ (IV.81)
Combining (IV.77)–(LABEL:c0dis) and (LABEL:dpitri), we obtain
$\begin{aligned} &\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)-\big(d\pi\circ d(f\circ\phi_{j}^{-1})\big)_{z^{\prime}}(v)\big|\\[7.11317pt] &\leq(2\alpha)^{-1}C_{j}\Big(\|\nabla\pi\|_{C(\overline{W^{\prime}})}+\text{Lip}(\nabla\pi|_{\overline{W^{\prime}}})\,\|\nabla(f\circ\phi_{j}^{-1})\|_{L^{\infty}(\phi_{j}(V_{j}\cap\overline{U}))}\Big)\\[7.11317pt] &\leq 2^{-1}C_{j}.\end{aligned}$ (IV.82)
By (IV.76), (LABEL:dpi>), (LABEL:<c_j) we finally conclude that for any $z\in\phi_{j}(\operatorname{supp}(\xi_{j})\cap\overline{U})$, $0<\varepsilon<\varepsilon_{0}$, $v\in\mathbb{S}^{n-1}$, it holds
@@
The estimate (IV.73) is obtained since $\overline{U}\subset\bigcup_{j=1}^{s}\operatorname{supp}(\xi_{j})$. ∎
@@
$\Lambda^{-1}|X|_{\tilde{g}}\leq|d(\vec{\Phi}_{k})(X)|\leq\Lambda|X|_{\tilde{g}},\qquad\text{for all }X\in T\Sigma\,\text{ and all }\,k\in\mathbb{N}.$ (IV.83)
@@
Now it remains to show that $\alpha_{k}\rightarrow\alpha$ in $C^{0}(\Sigma)$. Let $p\in\Sigma$. Since $h_{k}^{\prime}\rightarrow h$ in $C^{\infty}(\Sigma)$, by a similar argument as in Theorem IV.17 (replacing $\vec{e}_{i,k}$ by orthonormal frames on $T\Sigma$, and $d\vec{e}_{i,k}$ by covariant derivatives of these vector fields with respect to $h^{\prime}_{k}$), there exists a neighborhood $U$ of $p$ and $C^{\infty}$ diffeomorphisms $\varphi_{k}$ from $U$ to $\varphi_{k}(U)\subset\mathbb{R}^{2}$ such that for any $k\in\mathbb{N}\cup\{\infty\}$, it holds that $\varphi_{k}(p)=0$, $D^{2}\Subset\varphi_{k}(U)$, and $\varphi_{k}^{*}g_{\text{std}}$ is conformal to $h_{k}^{\prime}$ (write $h_{\infty}^{\prime}\coloneqq h$). Moreover, we have $\varphi_{k}\rightarrow\varphi_{\infty}$ in $C^{\infty}_{\textup{loc}}(U)$. Since $h_{k}^{\prime}$ is conformal to $g_{k}^{\prime}$, we also obtain that $\vec{\Phi}_{k}^{\prime}\circ\varphi_{k}^{-1}$ is conformal. We denote
@@
Then we have $\tilde{\lambda}_{k}\rightarrow\tilde{\lambda}_{\infty}$ in $C^{\infty}(\overline{D^{2}})$ since it holds $\varphi_{k}^{-1}\rightarrow\varphi_{\infty}^{-1}$ in $C^{\infty}(\overline{D^{2}})$ and $h_{k}^{\prime}\rightarrow h_{\infty}^{\prime}=h$ in $C^{\infty}(\Sigma)$. Since the sequence $\{\lambda_{k}\}$ is bounded in $L^{\infty}(D^{2})$ by (IV.83) and Theorem IV.19 (v), and $\vec{\Phi}_{k}^{\prime}\circ\varphi_{k}^{-1}\rightarrow\vec{\Phi}\circ\varphi_{\infty}^{-1}$ in $W^{2,2}(D^{2},\mathbb{R}^{m})$, we have $\lambda_{k}\rightarrow\lambda_{\infty}$ in $W^{1,2}(D^{2})$. Now by applying the argument in Corollary III.11 to $\{\vec{\Phi}_{k}^{\prime}\circ\varphi_{k}^{-1}\}$, we have $\lambda_{k}\rightarrow\lambda_{\infty}$ in $W^{2,1}_{\textup{loc}}(D^{2})$. Since $\alpha_{k}=(\lambda_{k}-\tilde{\lambda}_{k})\circ\varphi_{k}$ on $U$, and $p$ is arbitrarily chosen on $\Sigma$, we conclude that $\alpha_{k}\rightarrow\alpha$ in $W^{2,1}(\Sigma)$, hence also in $C^{0}(\Sigma)$ (see [81, Theorem 3.3.4 & 3.3.10]). ∎
@@ V Willmore surfaces
Motivated by the generalization of Willmore surfaces in higher dimension, we provide in this section a new proof of the regularity of Willmore surfaces which does not involve the choice of conformal coordinates. In Section V.1, we write the Euler–Lagrange equation of Willmore surfaces in divergence form. In Section V.2, we prove the regularity of Willmore surfaces.
@@ V.1 The Euler–Lagrange equation and conservation laws
Theorem V.1 (Noether theorem [157]).
@@
Thanks to the fact that $\Delta_{g}\vec{\Phi}\perp\partial_{l}\vec{\Phi}$, we have
@@
Combining (V.4)–(V.7), we obtain the pointwise a.e. variation
@@
In particular, it holds that $d\vec{L}\in W^{-1,2}+L^{1}(D^{2})$, and by Corollary III.5, we have $\vec{L}\in L^{2}_{\text{loc}}(D^{2})$. The quantity $\vec{L}$ alone is not sufficient to deduce some additional regularity since $\vec{L}$ does not possess additional derivatives a priori. Using the pointwise invariance of $H^{2}\,d\textup{vol}_{g}$ by dilations and rotations, we deduce some conservation laws related to $\vec{L}$ and $\vec{\Phi}$.
@@
Theorem V.4 ([171, Theorem 5.59]).
Let $\vec{\Phi}\in W^{2,2}_{\text{imm}}(D^{2},\mathbb{R}^{3})$ be a weak Willmore immersion. Then there exists $\vec{L}\in L^{2}_{\text{loc}}(D^{2},\mathbb{R}^{3})$ satisfying (V.15) [^10] Up to additive constants, this $\vec{L}$ is 1/2-1/2 of the one defined in [171, Theorem 5.59]. There is a typo in [171, Equation (5.214b)]: the ++ sign should be changed to -. and moreover, we have
@@
By differentiating the relations (V.19), we obtain a div-curl system in $\vec{R}$ and $S$.
Theorem V.6 ([171, Corollary 5.61]).
@@
Although $H^{2}\,d\textup{vol}_{g}$ is not pointwise invariant by inversions, this is the case of the Lagrangian $(H^{2}-K)\,d\textup{vol}_{g}$, see for instance [34]. In this section, we derive the associated conservation law (V.28), which is equivalent to (V.20c).
@@ V.2 Regularity of Willmore surfaces without conformal coordinates
By dilation and translation, it suffices to prove $\nabla S,\nabla\vec{R},\vec{H}\in M^{2,\gamma}(D_{1/2}(0))$. By definition of a weak immersion, the coefficients $a^{ij}\coloneqq g^{ij}(\det g)^{1/2}$ satisfy the uniform ellipticity condition (III.13). Fix a small positive number $\varepsilon_{0}$ which will be determined later. Since $\vec{\Phi}\in W^{2,2}(D^{2})$, we have $\nabla\vec{n}\in L^{2}(D^{2})$, hence there exists $r_{0}\in(0,\frac{1}{4})$ depending on $\epsilon_{0}$ and $\vec{\Phi}$ such that
@@
Let $p\in D_{1/2}(0)$, $r\in(0,r_{0})$. By Lemma III.3, we can define $\Psi_{S}\in W^{1,2}_{0}(D_{r}(p))$ to be the unique solution of (see Section III.1)
@@
We also define $\vec{\Psi}_{\vec{R}}\in W^{1,2}_{0}(D_{r}(p))$ to be the unique solution of
@@
Fix a small positive number $\varepsilon_{0}$ which will be determined later. As in the proof of Theorem V.8, there exists $r_{0}\in(0,\frac{1}{4})$ depending on $\epsilon_{0}$ and $\vec{\Phi}$ such that
@@
Furthermore, by Riesz representation theorem for Hilbert spaces, there exists a unique $\vec{\Psi}_{2}\in W^{1,2}_{0}(D_{r}(p))$ satisfying the estimate $\|\nabla\vec{\Psi}_{2}\|_{L^{2}(D_{r}(p))}\leq C(\Lambda)\,\|H\|_{L^{2}(D_{r}(p))}$ and the following equation holds
@@
The Hölder continuity of $\vec{n}$ then follows from standard knowledge of the Morrey–Campanato spaces (see for instance [153, Theorem 3.5.2] and [2]). ∎
@@
Let $\Sigma$ be a $2$-dimensional closed smooth manifold, $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(\Sigma,\mathbb{R}^{3})$. Then $\vec{\Phi}$ is a graph near $\vec{\Phi}(p)$ for any $p\in\Sigma$, and $\vec{\Phi}\in C^{\infty}$ under the graph coordinates.
@@
By [148, Eq. (1.5)] and standard estimates on the Riesz potential [1, Proposition 3.2(ii)], we obtain $M^{1,\tau}(D^{2})\hookrightarrow W^{-1,p}(D^{2})$ for any $p\in(2,\frac{2-\tau}{1-\tau})$. Now since $_{g}\big(d\vec{n}\mathbin{\dot{\wedge}}d\vec{R}\big)\in M^{1,\tau}_{\textup{loc}}(D^{2})$, we have $_{g}\big(d\vec{n}\mathbin{\dot{\wedge}}d\vec{R}\big)\in W^{-1,p}_{\textup{loc}}(D^{2})$ for some $p>2$. Applying [31, Theorem 3.1] to the equation (V.43a) then yields $\nabla S\in L^{p}_{\textup{loc}}(D^{2})$. Similarly, we have $\nabla\vec{R}\in L^{p}_{\textup{loc}}(D^{2})$, and hence by (V.43c), $H\in L^{p}_{\textup{loc}}(D^{2})$. Moreover, since $d(2H*_{g}d\vec{\Phi})\in W^{-1,p}_{\textup{loc}}(D^{2})$, combining (V.43d) and (V.44) with a similar argument as above implies $\nabla\vec{n}\in L^{p}_{\textup{loc}}(D^{2})$. It follows that
@@
Now return to the equations (V.43a)–(V.43d). The right-hand side of each equation is in $L^{q}_{\textup{loc}}(D^{2})$ for any $q<\infty$, and since $S,\vec{R},\vec{n},g\in W^{1,q}_{\textup{loc}}(D^{2})$, by [38, Theorem 4.1] and the proof of [66, Theorem 8.8], we obtain $S,\vec{R},\vec{n}\in W^{2,q}_{\textup{loc}}(D^{2})$, and then $\vec{\Phi}\in W^{3,q}_{\textup{loc}}(D^{2})$… We finally get $\vec{\Phi}\in C^{\infty}$ under the graph coordinates. ∎